2 |x+1| -2 x =|2 x -1|+1

2|x+1|-2x=|2x-1|+1

Grade:Upto college level

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear akash

2|x+1|-2x=|2x-1|+1

case 1   x <-1

2-(x+1)-2x=1-2x+1

so -(x+1) =1

 x=-2   it is a solution

case 2  -1≤x<0

so  

2(x+1)-2x=-2x+1+1

 2(x+1) =2

 x+1=1

 x=0     not a solution

case x≥0

so

2(x+1)-2x=2x-1+1

  2(x+1) =2(x+1)   this is true for any vvalue of x

so final solution is x=-2 and all value of x≥0

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin